In a hypothetical population which is in Hardy Weinberg equilibrium, the frequency for a recessive allele is 30%. What percentage of the population would be expected to show the dominant trait in the next generation?


Humans who are born homozygous for the recessive sickle cell allele die of sickle cell anemia, while those who are heterozygous are resistant to malaria (see chapter 4 for more information on this balanced polymorphism).  4% of the population of the Congo are homozygous recessive for the sickle cell allele.  What percentage of the population is resistant to malaria?


Red hair is a result of a recessive mutation at the melanocortin 1 receptor (MC1R).  As you learned, about 13% of the Scottish population is red haired. If the population mates randomly, what is the frequency of heterozygotes?


In the Ireland, 10% of the population is red haired.  Assuming mating is random, what is the frequency of heterozygotes in Ireland?
Answer #1

ANS: According the question frequency of recessive allele =30%

                                                                                              =30/100

                                                                                              = 0.3

hardy Weinberg equilibrium represented in ,

                     p2 + 2pq +q2 = 1

p =frequency of percentage

p + q = 1

q/p = represent the frequency of dominant or recessive allele

p = 0.3

p + q =1

q = 1- p

q = 1 - 0.3

q = 0.7

according to the question , the value of

p = 0.3

q = 0.7

substitute the value in the equation

p2 +2pq +q2

.09 +2(0.3) (0.7) +0.49

here only ask about dominant traits is q, so p2 neglected

0.42 + 0.49 = 0.91

91/100 X 100 = 91%

q =0.7